# Extended-Krylov-subspace methods for trust-region and norm-regularization subproblems

### v2

@@ 6.1 Other elliptical norms
For the general problem defined in terms of an elliptical norm $\|x\|_{S}=\sqrt{x^{T}Sx}$ for a given symmetric positive definite $S=P^{T}P$, a similar approach is possible. Transforming via (5.7) as in Subsection 5.2, the (non-interior) optimality condition (1.2) becomes
(6.3) $(A_{\mbox{\tiny P}}+\sigma I)x_{\mbox{\tiny P}}(\sigma)=b_{\mbox{\tiny P}}\;\;\mbox{with}\;\;\|x_{\mbox{\tiny P}}(\sigma)\|=\Delta$}
@@
$\sigma\geq\sigma_{\mbox{\tiny P}}{}_{\min}\stackrel{{\scriptstyle\rm def}}{{=}}\max(0,-\lambda_{1}(A_{\mbox{\tiny P}})),$}
@@
note that $\lambda_{1}(A_{\mbox{\tiny P}})$ is equivalently the leftmost eigenvalue $\lambda_{1}(A,S)$ of the symmetric matrix pencil $A+\lambda S$. The equivalent version of (6.1) is then
@@
$\theta=\sigma-\sigma_{\mbox{\tiny S}}\geq\sigma_{\mbox{\tiny P}}{}_{\min}-\sigma_{\mbox{\tiny S}}=\max(-\sigma_{\mbox{\tiny S}},-\lambda_{1}(A,S)-\sigma_{\mbox{\tiny S}}).$}
@@ 7 Numerical experiments
We now provide some evidence that our new approach is a useful alternative to common existing methods. We consider the set of 96 unconstrained test problems in the current release (2025-09-09) of the CUTEst optimization test examples [GoulOrbaToin15:coap] that have 1000 or more variables. For these, we generate the gradient and Hessian at the provided initial point, and use these for b-b and $A$ respectively.
We compare our new method (Algorithm 2 implemented as TREK), using the initial (Gershgorin-based) shifting strategy outlined in Section 6, with the multi-factorization method [GoulRobiThor10:mpc] available as TRS within GALAHAD [GoulOrbaToin03:toms] and the iterative method GLTR [GoulLuciRomaToin99:siopt] from the same library; GLTR only requires Hessian-vector products which can be an asset in some cases. Since all three methods can take advantage of information gathered when solving a problem with one radius and are now faced with another with identical data but a smaller radius, we consider each problem with radii, $\Delta=1$ and then $0.5$. TREK uses an iteration bound $m=100$, but almost always [^8] In four instances, $m$ was too small, but increasing the value to $400$ cured the problem. terminates with far-fewer $k$. Default, and roughly equivalent stopping rules were applied; TREK terminates as soon as (5.4) or (5.5) is smaller than $10^{-10}$.
@@
We summarize our findings in Figure 7.1 using the performance profile [dolamore02:mp] in which relative times are ranked. The profile is built using the data given in Appendix C.
@@
We did not expect any of the methods tested to be an overall winner, and indeed this is the case. Clearly, the factorization-free method GLTR is frequently very effective, particularly when factorization is a significant overhead. But we were encouraged to see that TREK performs well in other cases, and the requirement of a single factorization often proves decisive when compared to the multi-factorization approach used by TRS. [^9] TRS also failed in three cases because the shifted matrix $A+\sigma I$ became too close to singularity to progress.
@@ 8 Regularization subproblems
for some $\sigma\geq\sigma_{\min}$ [CartGoulToin11:mp, NestPoly06:mp]. Thus the manifold of solutions is identical to that for the trust-region subproblem, but the one of interest satisfies (8.2) instead of (1.5); there is no interior-solution possibility in this case.
@@ 9 Conclusions
We have provided an alternative to currently popular methods for solving quadratic trust-region and regularization problems. This lies somewhere between multiple factorization methods that were first popularised by Moré and Sorensen [MoreSore83:sisc] and CG/Lanczos-based factorization-free methods such as GLTR [GoulLuciRomaToin99:siopt]. The novel features are firstly that it is shown that the manifold of all possible solutions lies on an (approximately) very low-rank manifold, and secondly that by extending the Krylov subspace to include products with $A^{-1}$ as well as $A$, this manifold may be computed very efficiently. This then adds a further tool to the arsenal of solvers that trust-region and cubic-regularization methods for general optimization methods rely on. Unsurprisingly, the new method is not always the best, but in some cases it is. It is also remarkable that empirically that the extended Krylov-subspace approach requires so few iterations to build a good approximation of the solution manifold; examination of the results in Appendix C illustrates only a few instances where more than 20 iterations are required. We provide analysis to explain why this should be so, although this may well be tightened in future.
Software that implements the algorithms will be available very shortly [^10] The TREK algorithm with its documentation is now available as a beta release, and once C, Python, Julia and Matlab interfaces have been finished, it will be simple to extend these to the norm regularisation case discussed in Section 8 as the NREK package. as part of the GALAHAD [GoulOrbaToin03:toms] fortran library, and there are interfaces to other popular languages such as C, Python and Julia, and tools such as Matlab. The codes are designed to solve not only a single instance, but a sequence in which the regularization is tightened, without recourse to expensive re-evaluations; the current subspace $V_{k}$ is extended if it is insufficient as the regularization changes. The new packages will be rolled out as optional subproblem solvers in other GALAHAD packages in due course.
@@ Appendix A Convergence analysis
The first $j$-th matrix products of the EKS algorithm (Algorithm 1) generate an orthogonal basis for the extended Krylov subspace ${\cal K}_{j}(A,b)$. A natural question is then, how closely do the elements of this subspace lie to components of the manifold $x(\sigma)$. That is to say, for a given $\sigma$, is there an entry in ${\cal K}_{j}(A,b)$ close to $x(\sigma)$, and if so, how close?
Consider the symmetric positive-definite matrix $A$ with smallest and largest eigenvalues $\lambda_{1}=\lambda_{1}(A)$ and $\lambda_{n}=\lambda_{n}(A)$, and consider the interval ${\cal W}=[\lambda_{1},\lambda_{n}]\subset\mathbb{R_{+}^{*}}$. In this section, we repeat for completeness, with minor variations, the theory developed by Knizhnerman and Simoncini [KnizSimo11:NUMMAT] to prove that the convergence of the extended-Krylov-subspace method to build an approximate basis for the solution manifold $\{(A+\sigma I)^{-1}b\}$, for all $\sigma\notin-{\cal W}-(-\delta_{1},\delta_{n})=-(\lambda_{1}-\delta_{1},\lambda_{n}+\delta_{n})$, and $\delta_{1},\delta_{n}>0$, is linear at worst. The proof relies heavily on 19th to mid 20th century complex analysis, and we refer the reader to [KnizSimo11:NUMMAT] for the details; the technical buildup may appear somewhat fearsome to the uninitiated.
@@
The special (matrix-valued) case for which $u=A$ and $\Psi(w)=zI$ leads directly to (A.5).
@@
The extended Krylov method to approximate a solution manifold
@@
Appendix B Modified norm trust region subproblem solver
Here we provide a detailed description of the modifications to the TREK algorithm (Algorithm 2) that are necessary when using the trust region norm for which $\|x\|_{S}^{2}=x^{T}Sx$, with $S=P^{T}P$, as discussed in Section 5.2. As we stated there, it suffices to apply Algorithm 2 with the matrix $A_{\mbox{\tiny P}}=P^{-T}AP^{-1}$ and vector $b_{\mbox{\tiny P}}=P^{-T}b$, and to recover the solution via $x_{*}=P^{-1}x_{\mbox{\tiny P}}{}_{*}$. Since $S^{-1}=P^{-1}P^{-T}$, this is formally as Algorithm 4 4.
@@ Appendix B Modified norm trust region subproblem solver
Output: $x_{*}={\mbox{(approx)}\begin{array}[t]{c}\mbox{arg min}\vskip-2.84526pt\\ \mbox{ $\scriptstyle x\in\mbox{\raisebox{-0.8pt}{\footnotesize I\thinspace R\thinspace}}^{n}$ }\end{array}\;}{\scriptstyle\frac{1}{2}}x^{T}Ax-b^{T}x\;\;\mbox{subject to}\;\;\|Px\|\leq\Delta.$
@@
$\delta_{0}:=\|P^{-T}b\|$, $v_{0}:=P^{-T}b/\delta_{0}$, $u:=w/\delta_{0}$;
@@
Under the transformation $w_{k}=P^{-1}v_{k}$ i.e., $v_{k}=Pw_{k}$, and $u=Pv$, i.e., $v=P^{-1}u$ (and then replacing $v_{k}\rightarrow w_{k}$ and $u\rightarrow v$ so that the resulting changes may be easily compared with Algorithm 2), and introducing temporary workspace vectors $z$ and $s$, we may rewrite Algorithm 4 as Algorithm 5 5.
@@
While it might appear that Algorithm 5 requires multiple products of $S$ with selected vectors, as needed to calculate the terms $v_{i}^{T}Su$ that define $\alpha_{i}$ and $\beta_{i}$, fortunately this is not the case. There are two ways to avoid this. In the first, we maintain appropriate vectors $q_{i}\stackrel{{\scriptstyle\rm def}}{{=}}Sv_{i}$. We record this as Algorithm 6 6.
@@
Appendix C Detailed numerical experiments
Here we summarize the results from our numerical experiments that lead to Figure 7.1. As we said, we consider the set of 96 unconstrained test problems in the current release (2025-09-09) of the CUTEst optimization test examples [GoulOrbaToin15:coap] that have 1000 or more variables, evaluate the gradient and Hessian to provide b-b and $A$, and for each a pair of radii $\Delta=1$ and then $0.5$. We apply the GALAHAD solvers TREK (Algorithm 2), TRS [GoulRobiThor10:mpc] and GLTR [GoulLuciRomaToin99:siopt] in each case; when $\Delta=0.5$, we use data accumulated from the $\Delta=1$ case to “warm start” the second solve as each solver provides for this.
In Table LABEL:table-a1 we report the name of each test problem, the size of the instance chosen, $n$ (many CUTEst examples have multiple dimensions), the radius value, $\Delta$ and the optimal objective function found. For TREK, we give the terminal iteration count, $k_{\mbox{\scriptsize stop}}$, which is the number of solves with $A$ required, and the clock time required. For TRS, we report the number of factorizations and clock time required, while for GLTR, we state the number of vector products involving $A$ and again the time. The symbol $\dagger$} indicates an instance that the iteration bound $m=100$ was reached, while and $\ddagger$} indicates that a factorization failed through singularity.
Table C.1: Subproblem solves for CUTEst examples: Complete results

### v3

@@ 6.1 Other elliptical norms
For the general problem defined in terms of an elliptical norm $\|x\|_{S}=\sqrt{x^{T}Sx}$ for a given symmetric positive definite $S=^{T}$, a similar approach is possible. Transforming via (5.7) as in Subsection 5.2, the (non-interior) optimality condition becomes
(6.3) $(A_{\mbox{\tiny}}+\sigma I)x_{\mbox{\tiny}}(\sigma)=b_{\mbox{\tiny}}\;\;\mbox{with}\;\;\|x_{\mbox{\tiny}}(\sigma)\|=\Delta$}
@@
$\sigma\geq\sigma_{\mbox{\tiny}}{}_{\min}\stackrel{{\scriptstyle\rm def}}{{=}}\max(0,-\lambda_{1}(A_{\mbox{\tiny}})),$}
@@
note that $\lambda_{1}(A_{\mbox{\tiny}})$ is equivalently the leftmost eigenvalue $\lambda_{1}(A,S)$ of the symmetric matrix pencil [GoluvanL96, §7,7]. $A-\lambda S$. The equivalent version of (6.1) is then
@@
$\theta=\sigma-\sigma_{\mbox{\tiny S}}\geq\sigma_{\mbox{\tiny}}{}_{\min}-\sigma_{\mbox{\tiny S}}=\max(-\sigma_{\mbox{\tiny S}},-\lambda_{1}(A,S)-\sigma_{\mbox{\tiny S}}).$}
@@ 7 Numerical experiments
We now provide some evidence that our new approach is a useful alternative to common existing methods. We consider the set of unconstrained test problems in the current release (2025-09-09) of the CUTEst optimization test examples [GoulOrbaToin15:coap] that have 1000 or more variables. For these, we generate the gradient and Hessian at the provided initial point, and use these for b-b and $A$ respectively.
We compare our new method (Algorithm 2 implemented as TREK), using the initial (Gershgorin-based) shifting strategy outlined in Section 6, with the multi-factorization method [GoulRobiThor10:mpc] available as TRS within GALAHAD [GoulOrbaToin03:toms] and the iterative method GLTR [GoulLuciRomaToin99:siopt] from the same library; GLTR only requires Hessian-vector products which can be an asset in some cases. Since all three methods can take advantage of information gathered when solving a problem with one radius and are now faced with another with identical data but a smaller radius, we consider each problem with TREK uses an iteration bound , but almost always [^8] In instances, $m$ was too small, but increasing the value to $400$ cured the problem. terminates with far-fewer $k$. Default, and roughly equivalent stopping rules were applied; TREK terminates as soon as (5.4) or (5.5) is smaller than $10^{-10}$.
@@
We summarize our findings in Figure 7.1 using the performance profile [dolamore02:mp] in which relative times are ranked. The profile is built using the data given in Appendix D.
@@
We did not expect any of the methods tested to be an overall winner, and indeed this is the case. Clearly, the factorization-free method GLTR is frequently very effective, particularly when factorization is a significant overhead . But we were encouraged to see that TREK performs well in other cases, and the requirement of a single factorization often proves decisive when compared to the multi-factorization approach used by TRS. [^9] TRS also failed in cases because the shifted matrix $A+\sigma I$ became too close to singularity to progress.
@@ 8 Regularization subproblems
for some $\sigma\geq\sigma_{\min}$ Thus the of solutions is identical to that for the trust-region subproblem, but the one of interest satisfies (8.2) instead of (1.5); there is no interior-solution possibility in this case.
@@ 9 Conclusions
We have provided an alternative to currently popular methods for solving quadratic trust-region and regularization problems. This lies somewhere between multiple factorization methods that were first popularised by Moré and Sorensen [MoreSore83:sisc] and CG/Lanczos-based factorization-free methods such as GLTR [GoulLuciRomaToin99:siopt]. The novel features are firstly that it is shown that the of all possible solutions lies in an (approximately) very , and secondly that by extending the Krylov subspace to include products with $A^{-1}$ as well as $A$, this may be computed very efficiently. This then adds a further tool to the arsenal of solvers that trust-region and cubic-regularization methods for general optimization methods rely on. Unsurprisingly, the new method is not always the best, but in some cases it is. It is also remarkable that empirically that the extended Krylov-subspace approach requires so few iterations to good approximation of the solution ; examination of the results in Appendix D illustrates only a few instances where more than 20 iterations are required. We provide analysis to explain why this should be so, although this may well be tightened in future.
Software that implements the algorithms is now available 101010 as part of the GALAHAD [GoulOrbaToin03:toms] fortran library, and there are interfaces to other popular languages such as C, Python and Julia, and tools such as Matlab. The codes are designed to solve not only a single instance, but a sequence in which the regularization is tightened, without recourse to expensive re-evaluations; the current subspace $V_{k}$ is extended if it is insufficient as the regularization changes. The new packages will be rolled out as optional subproblem solvers in other GALAHAD packages in due course.
Acknowledgement
@@ Appendix A Convergence analysis
The first $j$-th matrix products of the EKS algorithm (Algorithm 1) generate an orthogonal basis for the extended Krylov subspace ${\cal K}_{j}(A,b)$. A natural question is then, how closely do the elements of lie to . That is to say, for a given $\sigma$, is there an entry in ${\cal K}_{j}(A,b)$ close to $x(\sigma)$, and if so, how close?
Consider the symmetric positive-definite matrix $A$ with smallest and largest eigenvalues $\lambda_{1}=\lambda_{1}(A)$ and $\lambda_{n}=\lambda_{n}(A)$, and consider the interval ${\cal W}=[\lambda_{1},\lambda_{n}]\subset\mathbb{R_{+}^{*}}$. In this section, we repeat for completeness, with minor variations, the theory developed by Knizhnerman and Simoncini [KnizSimo11:NUMMAT] to prove that the convergence of the extended-Krylov-subspace method to build an approximate basis for the ${\cal S}(A,b)$, for all $\sigma\notin-{\cal W}-(-\delta_{1},\delta_{n})=-(\lambda_{1}-\delta_{1},\lambda_{n}+\delta_{n})$, and $\delta_{1},\delta_{n}>0$, is linear at worst. The proof relies heavily on 19th to mid 20th century complex analysis, and we refer the reader to [KnizSimo11:NUMMAT] for the details; the technical buildup may appear somewhat fearsome to the uninitiated.
@@
The special (matrix-valued) case for which $u=A$ and $\Psi(w)=zI$, leads directly to (A.5).
@@
The extended Krylov method to approximate a
@@
Appendix B Computing an orthogonal basis of ${\cal K}_{\ell}(A,b)$
Given the $n$ by $n$ symmetric matrix $A$, the $n$-vector $b$, we define a sequences of extended Krylov subspaces
${\cal K}_{2k}(A,b)={\rm span}\{A^{-k}b,\ldots,A^{k-1}b\}$}
and
${\cal K}_{2k+1}(A,b)={\rm span}\{A^{-k}b,\ldots,A^{k}b\}={\rm span}\{{\cal K}_{2k}(A,b),A^{k}b\}$}
where ${\cal K}_{0}(A,b)={\rm span}\{b\}$, and thus
${\cal K}_{2k}(A,b)={\rm span}\{{\cal K}_{2k-1}(A,b),A^{-k}b\}.$}
We build a sequence of orthogonal basis for these so that
${\cal K}_{2k}(A,b)={\rm span}\{v_{-k},\ldots,v_{k-1}\}\;\;\mbox{and}\;\;{\cal K}_{2k+1}(A,b)={\rm span}\{v_{-k},\ldots,v_{k}\}$}
with ${\cal K}_{0}(A,b)=\{b/\|b\|\}$. Note that
$\begin{array}[]{rl}{\cal K}_{2k-1}(A,b)\hskip-5.69054pt&={\rm span}\{A^{-(k-1)}b,\ldots,A^{k-1}b\}={\rm span}\{v_{-(k-1)},\ldots,v_{k-1}\},\\ {\cal K}_{2k-2}(A,b)\hskip-5.69054pt&={\rm span}\{A^{-(k-1)}b,\ldots,A^{k-2}b\}={\rm span}\{v_{-(k-1)},\ldots,v_{k-2}\}\\ \;\;\mbox{and}\;\;{\cal K}_{2k-3}(A,b)\hskip-5.69054pt&={\rm span}\{A^{-(k-2)}b,\ldots,A^{k-2}b\}={\rm span}\{v_{-(k-2)},\ldots,v_{k-2}\}\end{array}$}
We find $v_{j}$ by orthogonalizing $Av_{-j}$ with respect to columns of $V_{2j}$, in which case
(B.11) $Av_{-j}\in{\cal K}_{2j+1},$}
and $v_{-(j+1)}$ by orthogonalizing $A^{-1}v_{j}$ with respect to the columns of $V_{2j+1}$, so that
(B.12) $A^{-1}v_{j}\in V_{2j+2}.$}
By construction, it then follows that
(B.13) $\delta_{k}v_{k}=Av_{-k}-\sum_{j=-k}^{k-1}\alpha_{j}v_{j}=Av_{-k}-\alpha_{-k}v_{-k}-\alpha_{k-1}v_{k-1}-\!\!\!\sum_{j=-k+1}^{k-2}\alpha_{j}v_{j}$}
for some normalizing constant $\delta_{k}$, which implies that
$0=\delta_{k}v_{i}^{T}v_{-k}=v_{j}^{T}Av_{-k}-\alpha_{j},$}
i.e.
(B.14) $\alpha_{j}=v_{j}^{T}Av_{-k}=v_{-k}^{T}Av_{j}$}
for $j=-k,\ldots,k-1$. Consider first $j=-k+1,\ldots,-1$, and for these
$\alpha_{j}\equiv\alpha_{-i}=v_{-i}^{T}Av_{-k}=v_{-k}^{T}Av_{-i}$}
for $i=1,\ldots,k-1$. But then, it follows from (B.11) that
$Av_{-i}\in{\cal K}_{2i+1}(A,b)\subseteq{\cal K}_{2k-1}(A,b),$}
and thus $\alpha_{-i}=0$ since $v_{-k}$ is orthogonal to ${\cal K}_{2k-1}(A,b)={\rm span}\{v_{-(k-1)},\ldots,v_{k-1}\}$. By contrast, consider $j=0,\ldots,k-2$, and note that in this case $v_{j}\in{\cal K}_{2k-3}(A,b)$. But then
$Av_{j}\in{\rm span}\{A^{-(k-3)}b,\ldots,A^{k-1}b\}\subset{\rm span}\{A^{-(k-1)}b,\ldots,A^{k-1}b\}={\cal K}_{2k-1}(A,b)$}
and thus as before $\alpha_{j}=0$ since $v_{-k}$ is orthogonal to ${\rm span}\{v_{-(k-1)},\ldots,v_{k-1}\}$. Thus (B.13) simplifies to a short-term recurrence
(B.15) $\delta_{k}v_{k}=Av_{-k}-\alpha_{-k}v_{-k}-\alpha_{k-1}v_{k-1}.$}
Similarly, by construction, we have that
(B.16) $\delta_{-(k+1)}v_{-(k+1)}=A^{-1}v_{k}-\sum_{j=-k}^{k}\beta_{j}v_{j}=A^{-1}v_{k}-\beta_{-k}v_{-k}-\beta_{k}v_{k}-\!\!\!\sum_{j=-k+1}^{k-1}\beta_{j}v_{j}$}
for another normalizing constant $\delta_{-(k+1)}$, which implies that
$0=\delta_{-(k+1)}v_{i}^{T}v_{-(k+1)}=v_{j}^{T}A^{-1}v_{-k}-\beta_{j},$}
i.e.
(B.17) $\beta_{j}=v_{j}^{T}A^{-1}v_{k}=v_{k}^{T}A^{-1}v_{j}$}
for $j=-k,\ldots,k$. Consider first $j=0,\ldots,k-1$, and for these (B.12) implies that
$A^{-1}v_{j}\in{\cal K}_{2i+2}(A,b)\subseteq{\cal K}_{2k}(A,b).$}
But then $\beta_{j}=0$ as $v_{k}$ does not lie in ${\rm span}\{v_{-k},\ldots,v_{k-1}\}$. Whereas, if $j=-k+1,\ldots,-1$,
$\beta_{j}\equiv\beta_{-i}=v_{-i}^{T}A^{-1}v_{k}=v_{k}^{T}A^{-1}v_{-i}$}
for $i=1,\ldots,k-1$. Note that in this case $v_{-i}\in{\cal K}_{2i}(A,b)\subseteq{\cal K}_{2k-2}(A,b)$, and then it follows that
$A^{-1}v_{-i}\in{\rm span}\{A^{-k}b,\ldots,A^{k-3}b\}\subset{\rm span}\{A^{-k}b,\ldots,A^{k-1}b\}={\cal K}_{2k}(A,b)$}
Hence, once again, $\beta_{j}=0$ as $v_{k}$ is orthogonal to ${\rm span}\{v_{-k},\ldots,v_{k-1}\}$. Thus, as before, (B.16) reduces to a short-term recurrence
(B.18) $\delta_{-(k+1)}v_{-(k+1)}=A^{-1}v_{k}-\beta_{-k}v_{-k}-\beta_{k}v_{k}.$}
The identities (B.14), (B.15), (B.17) and (B.18) then lead to Algorithm 1.
Appendix C Modified norm trust region subproblem solver
Here we provide a detailed description of the modifications to the TREK algorithm (Algorithm 2) that are necessary when using the trust region norm for which $\|x\|_{S}^{2}=x^{T}Sx$, with $S=^{T}$, as discussed in Section 5.2. As we stated there, it suffices to apply Algorithm 2 with the matrix $A_{\mbox{\tiny}}=^{-T}A^{-1}$ and vector $b_{\mbox{\tiny}}=^{-T}b$, and to recover the solution via $x_{*}=^{-1}x_{\mbox{\tiny}}{}_{*}$. Since $S^{-1}=^{-1}{}^{-T}$, this is formal as Algorithm 4 4.
@@ Appendix C Modified norm trust region subproblem solver
$\delta_{0}:=\|^{-T}b\|$, $v_{0}:=^{-T}b/\delta_{0}$, $u:=w/\delta_{0}$;
@@
Under the transformation $w_{k}=^{-1}v_{k}$ i.e., $v_{k}=w_{k}$, and $u=v$, i.e., $v=^{-1}u$ (and then replacing $v_{k}\rightarrow w_{k}$ and $u\rightarrow v$ so that the resulting changes may be easily compared with Algorithm 2), and introducing temporary workspace vectors $z$ and $s$, we may rewrite Algorithm 4 as Algorithm 5 5.
@@
While it might appear that Algorithm 5 requires multiple products of $S$ with selected vectors, as needed to calculate the terms $v_{i}^{T}Su$ that define $\alpha_{i}$ and $\beta_{i}$, fortunately this is not the case. There are two ways to avoid this. In the first, we maintain appropriate vectors $q_{i}\stackrel{{\scriptstyle\rm def}}{{=}}Sv_{i}$. We this as Algorithm 6 6.
@@
Appendix D Detailed numerical experiments
Here we summarize the results from our numerical experiments that lead to Figure 7.1. As we said, we consider the set of unconstrained test problems in the current release (2025-09-09) of the CUTEst optimization test examples [GoulOrbaToin15:coap] that have 1000 or more variables, evaluate the gradient and Hessian to provide b-b and $A$, We apply the GALAHAD solvers TREK (Algorithm 2), TRS [GoulRobiThor10:mpc] and GLTR [GoulLuciRomaToin99:siopt] in each case
In Table LABEL:table-a1 we report the name of each test problem, the size of the instance chosen, $n$ (many CUTEst examples have multiple dimensions), the radius value, $\Delta$ and the optimal objective function found. For TREK, we give the terminal iteration count, $k_{\mbox{\scriptsize stop}}$, which is the number of solves with $A$ required, and the clock time required. For TRS, we report the number of factorizations and clock time required, while for GLTR, we state the number of vector products involving $A$ and again the time. Both of the $\Delta$ regimes (a) and (b) are included for each solver. The symbol $\dagger$} indicates an instance that the iteration bound was reached, while and $\ddagger$} indicates that a factorization failed through singularity.
Table D.1: Subproblem solves for CUTEst examples: Complete results